OpenStudy (anonymous):

Can someone check my work on this please? Two forces with magnitudes of 200 and 100 pounds act on an object at angles of 60° and 170° respectively. Find the direction and magnitude of the resultant force. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (anonymous):

This is what I put: For the first triangle, u, I got u=<-98.48, 17.36>. For the second triangle, v, I got v=<100, 173.2>. Resultant force=u+v=<1.52,190.56>. Magnitude=~190.57. Direction angle= tan^1(b/a)=tan^1(190.56/1.52)=125.37.

OpenStudy (mathmale):

At first glance it appears as though you are on precisely the right track. I'm not actually checking the math, only the method you've used. As far as method: you're demonstrating good understanding of this material. Strongly suggest that you draw a diagram next time you try solving this kind of problem.

OpenStudy (anonymous):

I did draw the diagram, I just didn't draw it again on here. Thanks! :D God bless you!

OpenStudy (mathmale):

Always happy to hear from you and to be of help.